OpenStudy (anonymous):

If the point (5/13,12,13) corresponds to an angle theta in the unit circle, what is cot theta

jimthompson5910 (jim_thompson5910):

I think you meant to type (5/13, 12/13) right?

OpenStudy (anonymous):

yeah, my b

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

for any point (x,y) on the unit circle we can say tan(theta) = y/x and cot(theta) = x/y Note: tan is only defined when x is not zero, cot is only defined when y is not zero